re-write the equation as a function of x.
f(x)
-3x+4y=11

Answer by tutorcecilia(2144)

-3x+4y=11 [solve for the y-term; answer will be in the "y=mx+b" format]
-3x+3x+4y=-3x+11
4y=-3x+11
4x/4=-3x/4+11/4
y= -3/4x+11/4 [write as a function of x]
f(x)= -3/4x+11/4